OpenStudy (lucaz):

lim x->0 of (1+x)^(1/x)

OpenStudy (lucaz):

the answer is e, but how do I get there?

OpenStudy (anonymous):

one of two ways the first way, which is probably in the book is to "take the log" you get \[\ln(1+x)^{\frac{1}{x}}=\frac{1}{x}\ln(1+x)\] take the limit of that

OpenStudy (lucaz):

yeah, this limit is 1

OpenStudy (anonymous):

you get \(1\) and so the answer to the original problem is \(e^1\) that is one way

OpenStudy (anonymous):

the other way is to recognize one definition of \(e\) as \[\lim_{n\to \infty}(1+\frac{1}{n})^n\] or equivalently \[\lim_{x\to 0}(1+x)^{\frac{1}{x}}\]

OpenStudy (anonymous):

third way to to write what \[(1+x)^{\frac{1}{x}}\] really means, namely \[\large e^{\frac{1}{x}\ln(1+x)}\] and take the limit there it is the same work as the first way

OpenStudy (lucaz):

ok, with the first case, my textbook says ln y = ln f(x), then find the lim is 1, then.. lim of ln y = 1 I dont get the next step e^1, what happend to the lim of ln y ?

OpenStudy (anonymous):

you took the log as the first step

OpenStudy (anonymous):

get the limit is one but that is the limit of the log of the function , not the limit of the original function you have to go back to the original function \[\lim_{x\to a}\ln(f(x))=L\implies \lim_{x\to a}f(x)=e^L\]

OpenStudy (lucaz):

alright, thank you very much!